OpenStudy (anonymous):

actually all I really need to know is how to find (x+h)^1.5

OpenStudy (anonymous):

what do you actually actually need to do because there is nothing to "find" here. it is what it is

OpenStudy (anonymous):

you can rewrite it as \[(x+h)^{\frac{3}{2}}\] or as \[\sqrt{(x+h)^3}\] or even as \[(x+h)\sqrt{x+h}\] if you like

OpenStudy (anonymous):

I am trying to find the derivative of f(x) = (6x)^1.5 - (4x)^.5

OpenStudy (anonymous):

what you actually need is (x+h)^1.5= x+1.5*h*x^0.5+...O(h²)

OpenStudy (anonymous):

then it is \[9\sqrt{6x}-\frac{1}{\sqrt{x}}\]

OpenStudy (anonymous):

and you see, I have the answer in the back of my book, which is (9x)^.5 - (2x) ^(-.5)

OpenStudy (anonymous):

you want to use the limit method... dont you?

OpenStudy (anonymous):

I just really don't understand how to get there

OpenStudy (anonymous):

he didn't specify which method to use?

OpenStudy (anonymous):

the way path of solving it using the lim h...0 (f(x+h)-f(x)/)h is horrible if you havent seen binomial expansions or infinite series, I can tell you the "trick" if you want

OpenStudy (anonymous):

i will tell you what they used. they used the "power rule"

OpenStudy (anonymous):

oh yes, okay the limit method..duh. because we are finding it as the limit is approaching 0. yes. that is what you meant by lim h....0

OpenStudy (anonymous):

okay, what is the power rule?

OpenStudy (anonymous):

the derivative of x^n= n*x^(n-1)

OpenStudy (anonymous):

first off i am fairly sure that answer is wrong. lets check

OpenStudy (anonymous):

\[(6x)^{\frac{3}{2}}=\sqrt{6^3}x^{\frac{3}{2}}=6\sqrt{6}x^{\frac{3}{2}}\]

OpenStudy (anonymous):

so the derivative is \[\frac{3}{2}\times 6\sqrt{6}x^{\frac{1}{2}}=9\sqrt{6}\sqrt{x}=9\sqrt{6x}\]

OpenStudy (anonymous):

which is not the same as \[\sqrt{9x}\]

OpenStudy (anonymous):

second part is wrong too. what book is this?

OpenStudy (anonymous):

Lial Hunderford Holcomb Mathematics with applications 9th ed

OpenStudy (anonymous):

anyway , why would a book give an answer in this form: \[\sqrt{9x}\] and not \[3\sqrt{x}\]

OpenStudy (anonymous):

hmmm i have seen it. lial and hunferford used to be lial miller. @fiddlearound good point, and it is wrong in any case

OpenStudy (anonymous):

u sure it was (4x)^0.5 and not 4x^(0.5)

OpenStudy (anonymous):

@scountz it is a dumb question with decimal exponents and the answer given is wrong, so forget about it. or impress your teachers and say "i don't think this is right"

OpenStudy (anonymous):

oooh maybe it was \[6x^{\frac{3}{2}}+4x^{\frac{1}{2}}\]in which case the answer is \[9x^{\frac{1}{2}}+2x^{-\frac{1}{2}}\]

OpenStudy (anonymous):

aka \[9\sqrt{x}+\frac{2}{\sqrt{x}}\]

OpenStudy (anonymous):

but it is certainly not \[(9x)^{.5}\] for sure